# Generalised rank-constrained approximations of Hilbert–Schmidt operators on separable Hilbert spaces and applications

### v1

@@ 6 Application in Signal Processing, Reduced Rank Regression and Linear Operator Learning
In the field of signal processing, one common problem is to reconstruct a random input signal $x\in\mathbb{R}^{m}$ from a random observation signal $y\in\mathbb{R}^{m}$ that is related to $x$. One can assume that each observation satisfies $y=Gx+\varepsilon$ for some known linear map $G$ and additive noise $\varepsilon$, or weaker assumptions can be made on the relationship between $x$ and $y$. In [51], one considers the problem of reconstructing the signal $x$ as $Ay$, assuming additional structure only in the form of a known covariance matrix of $(x,y)$. To do so, the reconstruction quality is assessed in the expected squared reconstruction error, and thus the objective is to minimise $\mathbb{E}\lVert x-Ay\rVert^{2}$ over $A$. As the observation signal can be noisy, the rank of $A$ is restricted in order to filter out this noise. For more generality and motivated by applications, one can also allow $x$, $y$, and the reconstruction $Ay$ to be weighted by matrices $W_{x},W_{y}$ and $W_{A}$. The problem then becomes to find the rank- constrained minimiser $A$ of $\mathbb{E}\lVert W_{x}x-W_{A}AW_{y}y\rVert^{2}$. In [54], one considers an extension of this problem in which not just one signal but sets of signals $\{x(\alpha)\in\mathbb{R}^{m},\ \alpha\in K\}$ and $\{y(\alpha)\in\mathbb{R}^{m},\ \alpha\in K\}$ are considered, where $K$ is a suitable index set that may be countably or uncountably infinite, e.g. $K=\mathbb{N}$ or $K=[0,1]$, and $x(\alpha)$ and $y(\alpha)$ are finite-dimensional random vectors in $\mathbb{R}^{m}$ for every $\alpha$. Below, we consider the settings in [51, 54] in the case of infinite-dimensional signals, that is, in case the signals $x(\alpha)$ and $y(\alpha)$ are function-valued for every $\alpha\in K$.
@@
$\min\{\mathbb{E}\lVert W_{x}x-W_{A}AW_{y}y\rVert^{2},\ A\in\mathcal{B}_{00,r}(\mathcal{H}_{3},\mathcal{H}_{2})\}.$ (25)
Note that in infinite dimensions it is not a priori clear that the minimum in (10) is indeed attained for a bounded operator, c.f. Section 4, so 6.1 might not have a solution.
If ${F}=G=\mathbb{R}^{m}$, then we are back in the setting of [51]. If ${F}=G=L^{2}(K;\mathbb{R}^{m})$ with $(K,\mathcal{K},\lambda)$ a countably generated measure space with $\sigma$-finite measure $\lambda$, which is separable, then we are in the setting of [54], by interpreting the infinite set of signals $\{x(\alpha):\ \alpha\in K\}$ as one signal $x\in L^{2}(\Omega\times K;\mathbb{R}^{m})\simeq L^{2}(\Omega;L^{2}(K,\mathbb{R}^{m}))$. If $W_{x},W_{y}$ and $W_{A}$ are taken to be the identity, we are in the linear operator learning setting of [36] but with an additional rank restriction imposed on $A$. If $W_{x}=\Gamma=W_{A}$, $W_{y}=I$ and $x$ and $y$ are finite-dimensional and have mean 0, (25) reduces to the reduced rank setting of [27].
@@
We also define $C_{z}\coloneqq C_{zz}$. Hence with $x$ and $y$ as in 6.1, $C_{x}$ and $C_{y}$ are positive and self-adjoint, while $C_{xy}=C_{yx}^{*}$. We collect some further facts on covariances in Lemma 6.2, which is proven in Appendix B.
@@
Let $z\in L^{2}(\Omega;E_{1})$ and $w\in L^{2}(\Omega;E_{2})$. Then,
@@
    If $E_{1}=E_{2}$, then $\operatorname{tr}\,\left({C_{wz}}\right)=\mathbb{E}\langle w,z\rangle<\infty$ ,
@@
    $C_{Qw,z}=QC_{w,z}$ for $Q\in\mathcal{B}(E_{2},E_{3})$ and any separable Hilbert space $E_{3}$,
@@
    $C_{zw}=C_{w}^{1/2}UC_{z}^{1/2}$ for some $U\in\mathcal{B}(E_{1},E_{2})$ satisfying $\lVert U\rVert\leq 1$ and $U=P_{\overline{\operatorname{ran}{C_{w}^{1/2}}}}UP_{\overline{\operatorname{ran}{C_{z}^{1/2}}}}$.
Since $x,y$ in 6.1 are square integrable, we see from Lemma 6.2 item (i) that $C_{x}$, $C_{y}$ and $C_{xy}$ are all trace class. Thus $C_{x}^{1/2},C_{y}^{1/2}$ are Hilbert–Schmidt. We can now solve 6.1, assuming that $\operatorname{ran}{W_{A}}$ is closed. We note that in order to prove this result, we apply the adjoint formulation approach outlined in Remark 5.7.
@@
Suppose that $W_{A}$ has closed range. Then a solution to 6.1 exists if and only if there exists a truncated SVD $(P_{\overline{\operatorname{ran}{C_{y}^{1/2}W_{y}^{*}}}}(C_{y}^{1/2})^{\dagger}C_{yx}W_{x}^{*}P_{\operatorname{ran}{W_{A}}})_{r}$ satisfying
@@
In that case, a solution is given by
@@
Additionally, (27) uniquely specifies a solution of 6.1 that also satisfies (28) if and only if the truncated SVD in (27) is unique.
@@
Let $A\in\mathcal{B}_{00,r}(\mathcal{H}_{3},\mathcal{H}_{2})$. Let us also put $B\coloneqq C_{y}^{1/2}W_{y}^{*}\in\mathcal{B}(\mathcal{H}_{3},G),$ $C\coloneqq W_{A}^{*}\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2}),$ $N\coloneqq BA^{*}C$ and $M\coloneqq(C_{y}^{1/2})^{\dagger}C_{yx}W_{x}^{*}$. Then $\operatorname{rank}\left({N}\right)\leq\operatorname{rank}\left({A^{*}}\right)=\operatorname{rank}\left({A}\right)\leq r$, so $N$ is Hilbert–Schmidt. By Lemma 6.2 item (iii), $C_{yx}=C_{y}^{1/2}UC_{x}^{1/2}$ with $U\in\mathcal{B}(F,G)$ and $\operatorname{ran}{U}\subset\overline{\operatorname{ran}{C}_{y}^{1/2}}$. Hence, with (36) and Proposition A.3,
@@
Multiplying (29) by $C_{y}^{1/2}$ from the left and using (34), we see that $C_{y}^{1/2}(C_{y}^{1/2})^{\dagger}C_{yx}=C_{yx}.$ As all Hilbert spaces are over $\mathbb{R}$, it holds that $\operatorname{tr}\,\left({S^{*}T}\right)=\operatorname{tr}\,\left({ST^{*}}\right)$ for $S,T\in L_{2}$. Therefore,
@@
$=\lVert C_{y}^{1/2}W_{y}^{*}A^{*}W_{A}^{*}\rVert_{HS}^{2}+\lVert W_{x}C_{x}^{1/2}\rVert_{HS}^{2}$
@@
$=\lVert W_{x}C_{x}^{1/2}\rVert_{HS}^{2}+\lVert N\rVert_{HS}^{2}-2\operatorname{tr}\,\left({N^{*}M}\right)$
$=c+\lVert M-N\rVert_{HS}^{2}$
$=c+\lVert M-BA^{*}C\rVert_{HS}^{2}$ (30)
where $c\coloneqq\lVert W_{x}C_{x}^{1/2}\rVert_{HS}^{2}-\lVert M\rVert_{HS}^{2}$ does not depend on $A$. Since taking the adjoint is a bijective operation from $\mathcal{B}_{00,r}(\mathcal{H}_{3},\mathcal{H}_{2})$ onto $\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})$ and since $A^{**}=A$, we see from (30) that $A$ solves 6.1 if and only if $A^{*}$ solves
$\min\{\lVert M-BXC\rVert_{HS}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\}.$ (31)
Problem (31) does not have a solution when, with $\mathcal{X}_{r}$ as in Theorem 3.2, the problem
$\min\{\lVert M-BXC\rVert_{HS}:\ X\in\mathcal{X}_{r}\}.$ (32)
does not have a solution, since $B_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\subset\mathcal{X}_{r}$. If (32) does have a solution, then it has a solution $\hat{X}$ of the form (11) by Theorem 3.2, which then also solves (31). Indeed, by Proposition A.4, $\operatorname{ran}{W_{A}^{*}}$ is closed if and only if $\operatorname{ran}{W_{A}}$ is closed. By the assumption that $\operatorname{ran}{W_{A}}$ is closed and Proposition A.5 item (iv) it then follows that $(W_{A}^{*})^{\dagger}$ and hence $\hat{X}$ are bounded. Thus, 6.1, (31) and (32) are all equivalent, and $X$ solves 6.1 if and only if $X^{*}$ solves (31) and (32).
Using Theorem 3.2 we find that an optimal solution to (32) exists if and only if
@@
$P_{\ker{B}^{\perp}}\hat{X}P_{\overline{\operatorname{ran}{C}}}=\hat{X},$
@@
This solution is unique if and only if the rank-$r$ truncated SVD in its definition is unique. Since $\hat{X}^{*}$ solves 6.1, the proof is completed after taking the adjoint and applying Proposition A.5 item (v) and Proposition A.3 to $W_{A}^{*}$. ∎
Passing to the adjoint in the formulation (32) allows us to use Theorem 3.2, as in this case Theorem 3.2 gives bounded solutions, which are precisely the ones we search for in 6.1. If we would not pass to the adjoint formulation and instead try to use Theorem 3.2 on
$\mathbb{E}\lVert W_{x}x-W_{A}AW_{y}y\rVert^{2}=c+\lVert M^{*}-C^{*}AB^{*}\rVert_{HS}^{2}$
then since $B^{*}$ does not have closed range, we would in general find an unbounded minimiser $\tilde{A}$. Since 6.1 considers bounded solutions, such an unbounded minimiser cannot be a solution. While such a minimiser $\tilde{A}$ might satisfy $\lVert M^{*}-C^{*}\tilde{A}B^{*}\rVert_{HS}<\lVert M-B\hat{A}C\rVert_{HS}$, the mean square error $\mathbb{E}\lVert W_{x}x-W_{A}\tilde{A}W_{y}y\rVert^{2}$ now is ill-defined in general. Indeed, it requires $W_{y}y\in\operatorname{dom}{\tilde{A}}$ almost surely, which is a restrictive condition and is not satisfied in general. In particular, if $W_{y}=I$ and $y$ has a nondegenerate Gaussian distribution, then $\operatorname{dom}{\tilde{A}}=\operatorname{ran}{C_{y}^{1/2}}$ is the Cameron–Martin space of $y$ which has zero measure in infinite dimensions, see e.g. [5, Theorem 2.4.7]. Thus, $W_{y}y\not\in\operatorname{dom}{\tilde{A}}$ almost surely.
When no solution to 6.1 exists, it can still be approximated using the technique described in Section 3.2 and Remark 5.6, applied to the adjoint formulation (32).
@@
$\min\{\mathbb{E}\lVert x-Ay\rVert^{2},\ A\in\mathcal{B}_{00,r}(G,F)\}$ (33)
@@
We can apply Proposition 6.3 for the case $\mathcal{H}_{1}=\mathcal{H}_{2}=F$, $\mathcal{H}_{3}=G$ and $W_{A},W_{x},W_{y}$ equal to the identity operators. By (29) it holds that $\operatorname{ran}{(C_{y}^{1/2})^{\dagger}C_{yx}}\subset\overline{\operatorname{ran}{C_{y}^{1/2}}}$. The condition (26) therefore becomes $\operatorname{ran}{((C_{y}^{1/2})^{\dagger}C_{yx})_{r}}\subset\operatorname{ran}{C_{y}^{1/2}}$. Finally, (28) becomes $AP_{\ker{C_{y}^{1/2}}^{\perp}}=A$. It remains to show that $\ker{C_{y}^{1/2}}=\ker{C_{y}}$. If $0=Ah=A^{1/2}A^{1/2}h$, then $A^{1/2}h\in\ker{A^{1/2}}=\operatorname{ran}{A^{1/2}}^{\perp}$ using Proposition A.3. As also $A^{1/2}h\in\operatorname{ran}{A^{1/2}}$ it follows that $\ker{A}\subset\ker{A^{1/2}}$. The reverse inclusion is immediate. ∎
With $U$ as in the proof of Proposition 6.3, the existence condition $\operatorname{ran}{((C_{y}^{1/2})^{\dagger}C_{yx})_{r}}\subset\operatorname{ran}{C_{y}^{1/2}}$ is equivalent to $\operatorname{ran}{(UC_{x}^{1/2})_{r}}\subset\operatorname{ran}{C_{y}^{1/2}}$.
@@
We therefore see that the existence condition in [36] is stronger than the existence condition of Corollary 6.4. In particular, whenever one can find the optimal continuous linear operator $A$ without rank restriction, one can also find an optimal continuous and rank restricted operator. The converse does not hold in general.
@@
The minimality condition $\hat{A}=\hat{A}P_{\ker{C_{y}}^{\perp}}$ of Corollary 6.4 is equivalent to $\hat{A}(\ker{C}_{y})=0$. In [36, Proposition 3.11 and Remark 3.12(b)] it is shown that the solution of the unconstrained problem also satisfies this property, and in fact has maximal kernel among all solutions. The solution of the rank-constrained problem (33) can be seen to have maximal kernel among all solutions as well. Indeed, by applying Corollary 3.9 to (32) and taking the adjoint, we see that an arbitrary solution of (33) is of the form $\tilde{A}\coloneqq\hat{A}+T^{*}P_{\ker{C}_{y}}$ for some bounded $T:F\rightarrow G$. Now $\ker{T^{*}P_{\ker{C}_{y}}}^{\perp}\subset\ker{C}_{y}$. Hence $\ker{\tilde{A}}$ is maximal if and only if $T^{*}P_{\ker{C_{y}}}=0$, which holds if and only if $\tilde{A}=\hat{A}$. Furthermore, $\tilde{A}=\hat{A}$ holds if and only if $\hat{A}$ satisfies the minimality property $\hat{A}P_{\ker{C_{y}}^{\perp}}=\hat{A}$. Thus, the minimality property originating from (12) is exactly the maximal kernel property mentioned in [36, Proposition 3.11 and Remark 3.12(b)].

### v3

@@ 6 Application in Signal Processing, Reduced Rank Regression and Linear Operator Learning
In the field of signal processing, one common problem is to reconstruct a random input signal $x\in\mathbb{R}^{m}$ from a random observation signal $y\in\mathbb{R}^{m}$ that is related to $x$. One can assume that each observation satisfies $y=Gx+\varepsilon$ for some known linear map $G$ and additive noise $\varepsilon$, or weaker assumptions can be made on the relationship between $x$ and $y$. In [51], one considers the problem of reconstructing the signal $x$ as $Ay$, assuming additional structure only in the form of a known covariance matrix of $(x,y)$. To do so, the reconstruction quality is assessed in the expected squared reconstruction error, and thus the objective is to minimise $\mathbb{E}\lVert x-Ay\rVert^{2}$ over $A$. As the observation signal can be noisy, the rank of $A$ is restricted in order to filter out this noise. For more generality and motivated by applications, one can also allow $x$, $y$, and the reconstruction $Ay$ to be weighted by matrices $W_{x},W_{y}$ and $W_{A}$. The problem then becomes to find the rank-constrained minimiser $A$ of $\mathbb{E}\lVert W_{x}x-W_{A}AW_{y}y\rVert^{2}$. In [54], one considers an extension of this problem in which not just one signal but sets of signals $\{x(\alpha)\in\mathbb{R}^{m}:\ \alpha\in K\}$ and $\{y(\alpha)\in\mathbb{R}^{m}:\ \alpha\in K\}$ are considered, where $K$ is a suitable index set that may be countably or uncountably infinite, e.g. $K=\mathbb{N}$ or $K=[0,1]$, and $x(\alpha)$ and $y(\alpha)$ are finite-dimensional random vectors in $\mathbb{R}^{m}$ for every $\alpha$. Below, we consider the settings in [51, 54] in the case of infinite-dimensional signals, that is, in case the signals $x(\alpha)$ and $y(\alpha)$ are function-valued for every $\alpha\in K$.
@@
$\min\{\mathbb{E}\lVert W_{x}x-W_{A}AW_{y}y\rVert^{2}:\ A\in\mathcal{B}_{00,r}(\mathcal{H}_{3},\mathcal{H}_{2})\}.$ (27)
Note that in infinite dimensions it is not a priori clear that the minimum in (10) is indeed attained for a bounded operator, c.f. Section 4, so Problem 6.1 might not have a solution.
If ${F}=G=\mathbb{R}^{m}$, then we are back in the setting of [51]. If ${F}=G=L^{2}(K;\mathbb{R}^{m})$ with $(K,\mathcal{K},\lambda)$ a countably generated measure space with $\sigma$-finite measure $\lambda$, then $F$ and $G$ are separable, and we recover the setting of [54] by interpreting the infinite set of signals $\{x(\alpha):\ \alpha\in K\}$ as one signal $x\in L^{2}(\Omega\times K;\mathbb{R}^{m})\simeq L^{2}(\Omega;L^{2}(K,\mathbb{R}^{m}))$. If $W_{x},W_{y}$ and $W_{A}$ are taken to be the identity, we are in the linear operator learning setting of [36] but with an additional rank restriction imposed on $A$. If $W_{x}=\Gamma=W_{A}$, $W_{y}=I$ and $x$ and $y$ are finite-dimensional and have mean 0, then (27) reduces to the reduced rank setting of [27].
@@
We also define $C_{z}\coloneqq C_{zz}$. Hence with $x$ and $y$ as in Problem 6.1, $C_{x}$ and $C_{y}$ are positive and self-adjoint, while $C_{xy}=C_{yx}^{*}$. We collect some further facts on covariances in Lemma 6.2, which is proven in Appendix B.
@@
Let $z\in L^{2}(\Omega;E_{1})$ and $w\in L^{2}(\Omega;E_{2})$. Then the following hold:
@@
    If $E_{1}=E_{2}$, then $\operatorname{tr}\,\left({C_{wz}}\right)=\mathbb{E}\langle w,z\rangle<\infty$.
@@
    $C_{Qw,z}=QC_{w,z}$ for any separable Hilbert space $E_{3}$ and $Q\in\mathcal{B}(E_{2},E_{3})$.
@@
    $C_{wz}=C_{w}^{1/2}UC_{z}^{1/2}$ for some $U\in\mathcal{B}(E_{1},E_{2})$ satisfying $\lVert U\rVert\leq 1$ and $U=P_{\overline{\operatorname{ran}{C_{w}^{1/2}}}}UP_{\overline{\operatorname{ran}{C_{z}^{1/2}}}}$.
Since $x,y$ in Problem 6.1 are square integrable, we see from Lemma 6.2 item (i) that $C_{x}$, $C_{y}$ and $C_{xy}$ are all trace-class. Thus $C_{x}^{1/2},C_{y}^{1/2}$ are Hilbert–Schmidt. We can now solve Problem 6.1, assuming that $\operatorname{ran}{W_{A}}$ is closed. We note that in order to prove this result, we apply the adjoint formulation approach outlined in Remark 5.7.
@@
Suppose that $W_{A}$ has closed range. Then a solution to Problem 6.1 exists if and only if there exists a truncated SVD $(P_{\overline{\operatorname{ran}{C_{y}^{1/2}W_{y}^{*}}}}(C_{y}^{1/2})^{\dagger}C_{yx}W_{x}^{*}P_{\operatorname{ran}{W_{A}}})_{r}$ satisfying
@@
In this case, a solution is given by
@@
Additionally, (29) uniquely specifies a solution of Problem 6.1 that also satisfies (30) if and only if the truncated SVD in (29) is unique.
@@
Let $A\in\mathcal{B}_{00,r}(\mathcal{H}_{3},\mathcal{H}_{2})$. Let us also put $B\coloneqq C_{y}^{1/2}W_{y}^{*}\in\mathcal{B}(\mathcal{H}_{3},G),$ $C\coloneqq W_{A}^{*}\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2}),$ $N\coloneqq BA^{*}C$ and $M\coloneqq(C_{y}^{1/2})^{\dagger}C_{yx}W_{x}^{*}$. Then $\operatorname{rank}\left({N}\right)\leq\operatorname{rank}\left({A^{*}}\right)=\operatorname{rank}\left({A}\right)\leq r$, so $N$ is Hilbert–Schmidt. By Lemma 6.2 item (iii), $C_{yx}=C_{y}^{1/2}UC_{x}^{1/2}$ with $U\in\mathcal{B}(F,G)$ and $\operatorname{ran}{U}\subset\overline{\operatorname{ran}{C}_{y}^{1/2}}$. Hence, by (38) and Proposition A.3,
@@
Multiplying (31) by $C_{y}^{1/2}$ from the left and using $C_{yx}=C_{y}^{1/2}UC_{x}^{1/2}$, we see that $C_{y}^{1/2}(C_{y}^{1/2})^{\dagger}C_{yx}=C_{yx}.$ As all Hilbert spaces are over $\mathbb{R}$, it holds that $\operatorname{tr}\,\left({S^{*}T}\right)=\operatorname{tr}\,\left({ST^{*}}\right)$ for $S,T\in L_{2}$. Therefore,
@@
$=$ $\lVert C_{y}^{1/2}W_{y}^{*}A^{*}W_{A}^{*}\rVert_{L_{2}}^{2}+\lVert W_{x}C_{x}^{1/2}\rVert_{L_{2}}^{2}$
@@
$=$ $\lVert W_{x}C_{x}^{1/2}\rVert_{L_{2}}^{2}+\lVert N\rVert_{L_{2}}^{2}-2\operatorname{tr}\,\left({N^{*}M}\right)$
$=$ $c+\lVert M-N\rVert_{L_{2}}^{2}$
$=$ $c+\lVert M-BA^{*}C\rVert_{L_{2}}^{2}$ (32)
where $c\coloneqq\lVert W_{x}C_{x}^{1/2}\rVert_{L_{2}}^{2}-\lVert M\rVert_{L_{2}}^{2}$ does not depend on $A$. Since taking the adjoint is a bijective operation from $\mathcal{B}_{00,r}(\mathcal{H}_{3},\mathcal{H}_{2})$ onto $\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})$ and since $A^{**}=A$, we see from (32) that $A$ solves Problem 6.1 if and only if $A^{*}$ solves
$\min\{\lVert M-BXC\rVert_{L_{2}}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\}.$ (33)
Recall $\mathcal{X}_{r}$ from Setting 3.1, with $B$ and $C$ as defined above. The problem (33) does not have a solution when the problem
$\min\{\lVert M-BXC\rVert_{L_{2}}:\ X\in\mathcal{X}_{r}\}$ (34)
does not have a solution, since $B_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\subset\mathcal{X}_{r}$. If (34) does have a solution, then it has a solution $\hat{X}$ of the form (11) by Theorem 3.2, which then also solves (33). Indeed, by Proposition A.4, $\operatorname{ran}{W_{A}^{*}}$ is closed if and only if $\operatorname{ran}{W_{A}}$ is closed. By the assumption that $\operatorname{ran}{W_{A}}$ is closed and by Remark 3.5, $\hat{X}$ is bounded. Thus, Problem 6.1, (33) and (34) are all equivalent, and $X$ solves Problem 6.1 if and only if $X^{*}$ solves (33) and (34).
Recall the definitions of $B$ and $C$ at the start of the proof. Using Theorem 3.2 we find that an optimal solution to (34) exists if and only if
@@
$P_{\ker{B}^{\perp}}\hat{X}P_{\overline{\operatorname{ran}{C}}}=\hat{X}$
@@
This solution is unique if and only if the rank-$r$ truncated SVD in its definition is unique. Since $\hat{X}^{*}$ solves Problem 6.1, the proof is completed after taking the adjoint and applying Proposition A.5 item (v) and Proposition A.3 to $W_{A}^{*}$. ∎
Passing to the adjoint in the formulation (34) allows us to use Theorem 3.2, as in this case Theorem 3.2 gives bounded solutions, which are precisely the ones we search for in Problem 6.1. If we would not pass to the adjoint formulation and instead try to use Theorem 3.2 on
$\mathbb{E}\lVert W_{x}x-W_{A}AW_{y}y\rVert^{2}=c+\lVert M^{*}-C^{*}AB^{*}\rVert_{L_{2}}^{2},$
then since $B^{*}$ does not have closed range, we would in general find an unbounded minimiser $\tilde{A}$. Since Problem 6.1 considers bounded solutions, such an unbounded minimiser cannot be a solution. While such a minimiser $\tilde{A}$ might satisfy $\lVert M^{*}-C^{*}\tilde{A}B^{*}\rVert_{L_{2}}<\lVert M-B\hat{A}C\rVert_{L_{2}}$, the mean square error $\mathbb{E}\lVert W_{x}x-W_{A}\tilde{A}W_{y}y\rVert^{2}$ now is ill-defined in general. Indeed, it requires $W_{y}y\in\operatorname{dom}{\tilde{A}}$ almost surely, which is a restrictive condition and is not satisfied in general. In particular, if $W_{y}=I$ and $y$ has a nondegenerate Gaussian distribution, then $\operatorname{dom}{\tilde{A}}=\operatorname{ran}{C_{y}^{1/2}}$ is the Cameron–Martin space of $y$ which has zero measure in infinite dimensions, see e.g. [5, Theorem 2.4.7]. Thus, $W_{y}y\not\in\operatorname{dom}{\tilde{A}}$ almost surely.
When no solution to Problem 6.1 exists, it can still be approximated using the technique described in Section 3.2 and Remark 5.6, applied to the adjoint formulation (34).
@@
$\min\{\mathbb{E}\lVert x-Ay\rVert^{2}:\ A\in\mathcal{B}_{00,r}(G,F)\}$ (35)
@@
We can apply Proposition 6.3 for the case $\mathcal{H}_{1}=\mathcal{H}_{2}=F$, $\mathcal{H}_{3}=G$ and $W_{A},W_{x},W_{y}$ equal to the identity operators. By (31) it holds that $\operatorname{ran}{(C_{y}^{1/2})^{\dagger}C_{yx}}\subset\overline{\operatorname{ran}{C_{y}^{1/2}}}$. The condition (28) therefore becomes $\operatorname{ran}{((C_{y}^{1/2})^{\dagger}C_{yx})_{r}}\subset\operatorname{ran}{C_{y}^{1/2}}$. Finally, (30) becomes $AP_{\ker{C_{y}^{1/2}}^{\perp}}=A$. It remains to show that $\ker{C_{y}^{1/2}}=\ker{C_{y}}$. If $0=C_{y}h=C_{y}^{1/2}C_{y}^{1/2}h$, then $C_{y}^{1/2}h\in\ker{C_{y}^{1/2}}=\operatorname{ran}{C_{y}^{1/2}}^{\perp}$ using Proposition A.3. Since also $C_{y}^{1/2}h\in\operatorname{ran}{C_{y}^{1/2}}$, it follows that $\ker{C_{y}}\subset\ker{C_{y}^{1/2}}$. The reverse inclusion is immediate. ∎
With $U$ as in the proof of Proposition 6.3, the equation (31) implies that the existence condition $\operatorname{ran}{((C_{y}^{1/2})^{\dagger}C_{yx})_{r}}\subset\operatorname{ran}{C_{y}^{1/2}}$ is equivalent to $\operatorname{ran}{(UC_{x}^{1/2})_{r}}\subset\operatorname{ran}{C_{y}^{1/2}}$.
@@
We therefore see that the existence condition in [36] is stronger than the existence condition of Corollary 6.4. In particular, whenever one can find an optimal continuous linear operator $A$ that solves the problem in [36], i.e. (35) without rank restriction on $A$, then one can also find an optimal continuous and rank restricted operator that solves (35). The converse does not hold in general.
@@
The minimality condition $\hat{A}=\hat{A}P_{\ker{C_{y}}^{\perp}}$ of Corollary 6.4 is equivalent to $\hat{A}(\ker{C}_{y})=0$. In [36, Proposition 3.11 and Remark 3.12(b)] it is shown that the solution of the unconstrained problem also satisfies this property, and in fact has maximal kernel among all solutions. The solution of the rank-constrained problem (35) can be seen to have maximal kernel among all solutions as well. Indeed, by applying Corollary 3.9 to (34) and taking the adjoint, and using $\ker{C_{y}^{1/2}}=\ker{C_{y}}$ as explained in the proof of Corollary 6.4, we see that an arbitrary solution of (35) is of the form $\tilde{A}\coloneqq\hat{A}+T^{*}P_{\ker{C}_{y}}$ for some bounded $T:F\rightarrow G$. Now $\ker{(T^{*}P_{\ker{C}_{y}})}^{\perp}\subset\ker{C}_{y}$. Furthermore, $\ker{\hat{A}}^{\perp}\subset\ker{C_{y}}^{\perp}$ because $\hat{A}(\ker{\mathcal{C}_{y}})=0$. Hence $\ker{\tilde{A}}$ is maximal if and only if $T^{*}P_{\ker{C_{y}}}=0$, which holds if and only if $\tilde{A}=\hat{A}$. Furthermore, $\tilde{A}=\hat{A}$ holds if and only if $\tilde{A}$ satisfies the minimality property $\tilde{A}P_{\ker{C_{y}}^{\perp}}=\tilde{A}$. Thus, the minimality property originating from (12) is exactly the maximal kernel property mentioned in [36, Proposition 3.11 and Remark 3.12(b)].